find two consecutive integers whose product is 7 more than the square of the smaller number
Answer by Maths68(1474)

Let
Smaller integer = x
Greater integer= x+1
=======================================
Given
x*(x+1)=x^2+7
x*(x+1)=x^2+7
x^2+x=x^2+7
x^2-x^2+x=7
x=7
=====================================
Smaller integer = x =7
Greater = x+1 = 7+1=8
=========================================
Check
8*7=7^2+7
56=49+7
56=56